# Thomas–Fermi equation revisited: A variation on a theme by Majorana

### 31 Mar 2026 v1

@@ Abstract.
(Posted on the arXiv on 31 March 2026.)
@@ 1. Introduction
In 1927 and 1928, respectively, Thomas [9] and Fermi [7] managed to combine concepts of the old quantum theory (before mid-1925, that is) with the then-recent methods of quantum mechanics — which grew rapidly from the seminal works of Heisenberg, Schrödinger, and others — for a first successful model of many-electron atoms. The central mathematical object, the second-order differential equation now known as the Thomas–Fermi (TF) equation, immediately caught Majorana’s attention, who had the profound insight that there are scaling properties that can be exploited for establishing an equivalent first-order differential equation.
@@
In Sec. 2, we set the stage by recalling the second-order TF equation and introducing three scale-invariant functions that are constructed from the solutions of the TF equation. Next, in Sec. 3, we establish Majorana’s first-order equation by an argument that is closer in spirit to that in [1] than to the original one in Majorana’s notes, as recalled by Esposito [5]. The analogous argument for the second solution of the TF equation is presented in Sec. 4, where we report the other Majorana-type first-order equation.
Sections 5 and 6 deal with the power series for the two equations, respectively. The first series is the one that Esposito introduced [5] and the second is constructed in close analogy. Both series are used for computing anew the basic numbers associated with the TF functions, which are then compared with the numbers found in the 1980s. Further applications, in particular the computation of various relevant integrals involving the neutral-atom TF function, are the subject matter of Sec. 7. These are then used in Sec. 8 for calculating anew the coefficients in the formulas for the binding energy of neutral atoms and also the leading term in the expression for the ionization energy.
@@ 2. Setting the stage
with $Q^{3}=R^{3}P^{8}$. They are scale-invariant in the sense that the replacement $f\to f_{\lambda}$ yields $P(x)\to P(\lambda x)$, $Q(x)\to Q(\lambda x)$, and $R(x)\to R(\lambda x)$ without additional powers of $\lambda$ as overall factors.
@@ 3. Majorana’s equation for case (i)
(3.7) $U(t)=\int_{0}^{t}\mathrm{d}t^{\prime}\,\frac{t^{\prime}u(t^{\prime})}{1-{t^{\prime}}^{2}u(t^{\prime})}$
@@ 8. Energy formulas
In atomic units ($1\,\mathrm{Ha}=27.211\,386\,246\,\mathrm{eV}$), the binding energy of a neutral atom with $Z$ electrons is
@@ 9. Summary and outlook
(9.2) $f(0)=1\,,\quad f(x_{0})=0\,,\quad-x_{0}f^{\prime}(x_{0})=q\quad\mbox{for}\quad 0<q<1\,$,
where the value of the positive $x_{0}$ depends on $q$, the degree of ionization. [^7] The two cases in (2.3) and (2.4) are obtained in the limit of $q\to 0$}, $x_{0}\to\infty$} without and with the scaling for $\lambda=x_{0}$}, respectively. This is unexplored territory.

### 23 Apr 2026 v2

@@ Abstract.
(Posted on the arXiv on 23 April 2026.)
@@ 1. Introduction
In 1926 and 1927, respectively, Thomas [13] and Fermi [7] managed to combine concepts of the old quantum theory (before mid-1925, that is) with the then-recent methods of quantum mechanics — which grew rapidly from the seminal works of Heisenberg, Schrödinger, and others — for a first successful model of many-electron atoms. The central mathematical object, the second-order differential equation now known as the Thomas–Fermi (TF) equation, caught Majorana’s attention, who had the profound insight that there are scaling properties that can be exploited for establishing an equivalent first-order differential equation.
@@
We begin with a brief account of Thomas’s and Fermi’s reasoning in Sec. 2. Then, in Sec. 3, we set the stage by recalling the second-order TF equation and introducing three scale-invariant functions that are constructed from the solutions of the TF equation. Next, in Sec. 4, we establish Majorana’s first-order equation by an argument that is closer in spirit to that in [1] than to the original one in Majorana’s notes, as recalled by Esposito [5]. The analogous argument for the second solution of the TF equation is presented in Sec. 5, where we report the other Majorana-type first-order equation.
Sections 6 and 7 deal with the power series for the two equations, respectively. The first series is the one that Esposito introduced [5] and the second is constructed in close analogy. Both series are used for computing anew the basic numbers associated with the TF functions, which are then compared with the numbers found in the 1980s. Further applications, in particular the computation of various relevant integrals involving the neutral-atom TF function, are the subject matter of Sec. 8. These are then used in Sec. 9 for calculating anew the coefficients in the formulas for the binding energy of neutral atoms and also the leading term in the expression for the ionization energy.
@@
2. How Thomas and Fermi reasoned
Thomas and Fermi, independently, wondered about many-electron atoms, which were beyond the reach of the then-new quantum mechanics. The Schrödinger wave function for $N$ electrons has $2^{N}$ components, each a complex-valued function of the $3N$ electron coordinates. A compact analytical expression for the atomic ground-state wave function was out of the question then, as it is now, and there was no method for obtaining meaningful approximations. [^1] Hartree’s first step in this direction [9] was still in the future.
Thomas and Fermi resorted to modeling the atom by picturing the electrons moving independently under the influence of the force associated with an effective potential energy $V(\vecfont{r})$, a real-valued function of an electron’s position vector $\vecfont{r}$, which they approximated by
(2.1) $V(\vecfont{r})=-\frac{Ze^{2}}{r}+\frac{Ze^{2}}{r_{\mathrm{ch}}^{\ }}+e^{2}\int(\mathrm{d}\vecfont{r}^{\prime})\,\frac{n(\vecfont{r^{\prime}})}{\mathopen{\boldsymbol{|}}\vecfont{r}-\vecfont{r}^{\prime}\mathclose{\boldsymbol{|}}}\,$,
where $e$ is the elementary charge, $r=\mathopen{\boldsymbol{|}}\vecfont{r}\mathclose{\boldsymbol{|}}$} is the length of $\vecfont{r}$, $(\mathrm{d}\vecfont{r}^{\prime})$ denotes the euclidean volume element associated with $\vecfont{r}^{\prime}$, and $n(\vecfont{r}^{\prime})$ is the density of electrons at $\vecfont{r}^{\prime}$; the first term is the Coulomb potential of the nucleus with charge $Ze$, the second term is the chemical potential parameterized by the distance $r_{\mathrm{ch}}^{\ }$, and the third term is the electrostatic potential of the charge density $en(\vecfont{r})$ of the electrons. Note that, if the wave function were at hand, one would obtain $n(\vecfont{r})$ by integrating the sum of the absolute squares of all components over the coordinates of N1{N-1} electrons.
The second ingredient is the statistical-physics expression for the electron density in terms of the momentum integral over the region where the single-electron energy $\vecfont{p}^{2}/(2m)+V(\vecfont{r})$ is negative, [^2] As regulated by the chemical potential.
(2.4) $n(\vecfont{r})$ $=\frac{2}{(2\pi\hbar)^{3}}\int(\mathrm{d}\vecfont{p})\;\Biggl\{\begin{array}[]{cl}1&\displaystyle\text{where}\ \vecfont{p}^{2}+2mV(\vecfont{r})<0\\[2.15277pt] 0&\text{elsewhere}\end{array}\Biggr\}$
$=\frac{1}{3\pi^{2}\hbar^{3}}\Bigl(-2mV(\vecfont{r})\Bigr)_{+}^{3/2}\,$,
counting two electrons per phase-space volume $(2\pi\hbar)^{3}$; here, $m$ is the electron mass, $\vecfont{p}$ is the electron’s momentum vector, $2\pi\hbar$ is Planck’s constant, and $\bigl(y\bigr)_{+}=\bigl(y+\mathopen{\boldsymbol{|}}y\mathclose{\boldsymbol{|}}\bigr)/2$} suppresses negative values of the real variable $y$. The approximations entering (2.1) and (2.4) are semi-classical in nature; both can be refined by including quantum corrections, which have been studied extensively in the century since Thomas’s and Fermi’s seminal work.
Upon applying the Laplace differential operator on both sides of (2.1),
(2.5) $r>0\,:\quad-\boldsymbol{\nabla}^{2}V(\vecfont{r})=4\pi e^{2}n(\vecfont{r})\,$,
and taking into account that an isolated atom is isotropic by writing
(2.6) $V(\vecfont{r})=-\frac{Ze^{2}}{r}f(\kappa r)\quad\text{with}\quad\kappa\frac{\hbar^{2}}{me^{2}}=8(6\pi)^{-2/3}Z^{1/3}\,$,
we arrive at the familiar TF differential equation for $f(x)$ in (3.1) below. The combination of physical constants that multiplies $\kappa$ is Bohr’s radius, the natural unit of length in atomic physics, and the numerical factor is approximately $61/54$.
The first steps taken by Thomas and Fermi triggered a large body of work in the next two decades, summarized in Gombás’s classic of 1949 [8]. In the systematic development of density functional theory from the mid-1960s onward, the TF model is the basic approximation, and one has learned, in which sense the TF approximation is exact when both $N$ and $Z$ are very large. Recent accounts on the state of affairs from the physical and the mathematical perspectives are the chapters by Okun and Burke [11] and Siedentop [12], respectively, in the proceedings of a workshop in 2019.
@@ 3. Setting the stage
with $Q^{3}=R^{3}P^{8}$. They are scale-invariant in the sense that the replacement $f\to f_{\lambda}$ yields $P(x)\to P(\lambda x)$, $Q(x)\to Q(\lambda x)$, and $R(x)\to R(\lambda x)$ without additional powers of $\lambda$ as overall factors. In (3.6) and (3.8), f(x)>0{f(x)>0} is assumed; for $f(x)\leq 0$}, put P(x)=0{P(x)=0} and $R(x)^{-1}=0$}.
@@ 4. Majorana’s equation for case (i)
(4.7) $U(t)=\int_{0}^{t}\mathrm{d}t^{\prime}\,\frac{t^{\prime}u(t^{\prime})}{1-{t^{\prime}}^{2}u(t^{\prime})}\,$,
@@ 9. Energy formulas
In atomic units ($1\,\mathrm{Ha}=27.211\,386\,246\,\mathrm{eV}$), the binding energy of a neutral atom with N=Z{N=Z} electrons is
@@ 10. Summary and outlook
(10.2) $f(0)=1\,,\quad f(x_{\mathrm{ch}}^{\ })=0\,,\quad-x_{\mathrm{ch}}^{\ }f^{\prime}(x_{\mathrm{ch}}^{\ })=q\quad\mbox{for}\quad 0<q<1\,$,
where the value of the positive $x_{\mathrm{ch}}^{\ }$ depends on q=1N/Z{q=1-N/Z}, the degree of ionization. [^9] The relation $x_{\mathrm{ch}}^{\ }=q\kappa r_{\mathrm{ch}}^{\ }$} connects with (2.6). The two cases in (3.3) and (3.4) are obtained in the limit of $q\to 0$}, $x_{\mathrm{ch}}^{\ }\to\infty$} without and with the scaling for $\lambda=x_{\mathrm{ch}}^{\ }$}, respectively. This is unexplored territory.
@@
Acknowledgment
I am extremely grateful for the long-standing support from the Centre for Quantum Technologies, Singapore, where part of this work was done.